OpenStudy (radball4):

Which of the following sets is closed under multiplication? {0, 1, 2, 4} {0, 1} {1, 2} None of the sets are closed under multiplication.

OpenStudy (mathmate):

A set is closed under multiplication if we take any two numbers out of the set and multiply, the result is a member of the set. The two numbers may be repeated, i.e. multiply by itself. For example: X={0,-1} is not closed because 0*0=0 \(\in\) X, 0*(-1)=0 \(\in\) X, but (-1)*(-1)=1 \(\notin\) X, so X is not closed under multiplication. You can repeat the same test for all the given sets to get your answer.